OpenStudy (anonymous):

Free medal and Fan Find f(5) for this sequence: f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

Vocaloid (vocaloid):

first step is to find f(3) if f(n) = f(1) + f(2) + f(n-1) what is f(3)?

OpenStudy (anonymous):

5

Vocaloid (vocaloid):

not quite, please try again f(3) = f(1)+f(2)+f(2)

OpenStudy (anonymous):

4

Vocaloid (vocaloid):

one more time... f(1) = 2 f(2) = 3 f(3) = f(1) + f(2) + f(2) = ?

OpenStudy (anonymous):

@Vocaloid Im sorry i really dont know

Vocaloid (vocaloid):

 fill in the blank

OpenStudy (anonymous):

8 @Vocaloid

Vocaloid (vocaloid):

right, f(3) = 8 now we need to find f(4) f(1) = 2 f(2) = 3 f(3) = 8 f(4) = f(1) + f(2) + f(3) = __?

OpenStudy (anonymous):

13

Vocaloid (vocaloid):

great! we're almost done f(1) = 2 f(2) = 3 f(4) = 13 f(5) = f(1) + f(2) + f(4) = ?

OpenStudy (anonymous):

18

Vocaloid (vocaloid):

great, we're done

OpenStudy (anonymous):

Thanks for the help @Vocaloid